# Nonhermitian topological zero modes at smooth domain walls: Exact solutions

### v1

@@ III.3.1 One zero mode for $W_{R}=\pm 1$
For $\operatorname{sgn}({M}_{R})=-1$ and $\mathrm{Re}({v}_{R})\neq 0$, the topological invariant is nonzero $W_{R}=\operatorname{sgn}(\mathrm{Re}({v}_{R}))\neq 0$, giving $\operatorname{sgn}(\mathrm{Re}(\alpha_{R}^{\pm}))=s\operatorname{sgn}(\mathrm{Re}({v}_{R}))$. Choosing $s=-\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ gives $\mathrm{Re}(\alpha_{R}^{\pm})<0$, and therefore the two linearly independent solutions $\varphi_{1,2}^{s}(x)$ in Eq. 42 diverge for $x\to\infty$ with different decaying lengths, and thus their linear combination cannot satisfy the boundary conditions. Choosing instead $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ gives $\mathrm{Re}(\alpha_{R}^{\pm})>0$, and with both solutions $\varphi_{1,2}^{s}(x)$ decaying to zero for $x\to\infty$. Therefore, the particular solution satisfying the boundary conditions is the linear combination
@@
with pseudospin $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ in Eq. 2, $\varphi_{1,2}^{s}(x)$ as in Eqs. 42 and 57, ${\alpha_{L}}$, ${\alpha_{R}}$ given by Eq. 37, and with $A_{1,2}$ chosen such that $\varphi(0)=A_{1}\varphi_{1}^{s}(0)+A_{2}\varphi_{2}^{s}(0)=0$. Since $\mathrm{Re}(\alpha_{R}^{-})\leq\mathrm{Re}(\alpha_{R}^{+})$, the asymptotic behavior at large distances is given by
@@
where $\mu_{R}^{-}=\mathrm{Re}(\alpha_{R}^{-})=-{M}_{R}$. If one can expand ${m}(x)$ and ${v}(x)$ as in Eq. 46, taking $F_{1,2}$ as in Eq. 57, the constants $A_{1,2}$ must satisfy
@@
Notice that, when considering the solutions of the modified Jackiw-Rebbi equation in Eq. 3 on the interval $0\leq x<\infty$, the choice of the fields at $x=-\infty$ is somewhat arbitrary. Therefore, one can always choose the values of ${v}_{L}$ and ${m}_{L}$ in order to match the conditions ${v}_{L,R}={v}$, ${v}_{L,R}={v}$ and ${m}_{L,R}={m}$, ${v}_{L,R}=\pm{v}$, ${v}_{L,R}=\pm{v}$ and ${m}_{L,R}={m}$ in Table I of Ref. marra_topological_2024, . and obtain the respective simplified forms of the wavefunctions. However, if one can expand ${m}(x)$ and ${v}(x)$ as in Eq. 46, the functional form of the fields ${m}(x)$ and ${v}(x)$ on the interval $0\leq x<\infty$ determines univocally the values of the fields at $x=-\infty$.
The asymptotic behavior is determined by the exponent $\alpha_{R}^{-}=s{v}_{R}-{q}_{R}$ (characteristic lengths $\xi_{R}^{-}$, $\lambda_{R}^{-}$). This is because, even though the two solutions have different decaying lengths $\xi_{R}^{\pm}$, one has that $\xi_{R}^{-}\geq\xi_{R}^{+}$, which means that the decaying length $\xi_{R}^{-}$ dominates at large distances. In the hermitian case $\mathrm{Im}({m}(x))=\mathrm{Im}({v}(x))=0$, it is always possible to find a set of $A_{1,2}$ satisfying the equations above and such that $\varphi(x)$ is real.
@@ III.3.2 No zero modes otherwise
If $\operatorname{sgn}({M}_{R})\geq 0$ or if $\mathrm{Re}({v}_{R})=0$, no nontrivial linear combination of the general solutions can satisfy the boundary conditions. For $\operatorname{sgn}({M}_{R})=1$, the topological invariant is zero $W_{R}=0$, and one has $\mathrm{Re}(\alpha_{R}^{+})>0>\mathrm{Re}(\alpha_{R}^{-})$. Therefore only the solution $\varphi_{2}^{s}(x)$ in Eq. 42 decays to zero for $x\to\infty$. For $\operatorname{sgn}({M}_{R})=0$, one has that either $\alpha_{R}^{+}=0$ or $\alpha_{R}$, depending on the choice of the pseudospin. Therefore, only one of the solutions $\varphi_{1,2}^{s}(x)$ decays to zero for $x\to\infty$. However, since in general $\varphi_{1,2}^{s}(0)\neq 0$, the boundary conditions cannot be satisfied. For ${m}_{R}<0$ (or $\operatorname{sgn}({M}_{R})=-1$) and $\mathrm{Re}({v}_{R})=0$, one has $\mathrm{Re}(\alpha_{R}^{\pm})=0$. Hence, neither solutions $\varphi_{1,2}^{s}(x)$ decay to zero for $x\to\infty$, and consequently, the boundary conditions cannot be satisfied.
@@ IV.1 Oscillating behavior
Here, we will see that the quantity ${K}_{L,R}$ will discriminate between oscillatory and nonoscillatory behavior, generalizing to the nonhermitian case the analogous quantity introduced in Ref. marra_topological_2024, for hermitian systems. One has
$\begin{array}[]{ll}\mathrm{Im}(\alpha_{L}^{\pm})=0,&\text{ for }{K}_{L}=0,\\ \mathrm{Im}(\alpha_{L}^{\pm})\neq 0&\text{ for }{K}_{L}\neq 0,\end{array}$ (25)
@@
where we remind that $\mathrm{Re}(\alpha_{L}^{\pm})=\mu_{L}^{\pm}$ and $\mathrm{Im}(\alpha_{L}^{\pm})=\kappa_{L}^{\pm}$. The case where ${K}_{L}\neq 0$ corresponds to oscillating behavior for $x\to-\infty$, while ${K}_{L}=0$ corresponds to no oscillations. If ${m}_{L}\in\mathbb{R}$, we see that $\operatorname{sgn}({M}_{L})=\operatorname{sgn}({m}_{L})$, and thus the conditions in Eq. 12 simplify by substituting ${M}_{L}\to{m}_{L}$. Moreover, if the system is asymptotically hermitian for $x\to-\infty$, i.e., if ${m}_{L},{v}_{L}\in\mathbb{R}$, one has that ${K}_{L}=0$ if ${{v}}_{L}^{2}+{m}_{L}\geq 0$ and ${K}_{L}\neq 0$ otherwise. This gives
$\begin{array}[]{ll}\mathrm{Im}(\alpha_{L}^{\pm})=0,&\text{ for }{{v}}_{L}^{2}+{m}_{L}>0,\\ \mathrm{Im}(\alpha_{L}^{\pm})=0,\quad\mathrm{Re}(\alpha_{L}^{+})=\mathrm{Re}(\alpha_{L}^{-}),&\text{ for }{{v}}_{L}^{2}+{m}_{L}=0,\\ \mathrm{Im}(\alpha_{L}^{\pm})\neq 0,\quad\mathrm{Im}(\alpha_{L}^{+})=-\mathrm{Im}(\alpha_{L}^{-}),&\text{ for }{{v}}_{L}^{2}+{m}_{L}<0,\end{array}$ (27)
@@
The oscillatory behavior of the modes is determined by the quantities ${K}_{L,R}$. For ${K}_{L,R}=0$, one has $\kappa_{L,R}=0$, respectively, which mandates that the modes decay exponentially for $x\to\mp\infty$, respectively. Conversely, for ${K}_{L,R}\neq 0$, one has $\kappa_{L,R}\neq 0$, respectively, which mandates modes with exponentially damped oscillations for $x\to\mp\infty$, respectively. In summary, one has
@@
This is also summarized in Table 2.
@@ IV.2 Reality
Note also that, in the hermitian case $\mathrm{Im}({m}(x))=\mathrm{Im}({v}(x))=0$, the modified Jackiw-Rebbi equation is real, and consequently, all solutions are real for some choices of $A_{1,2}$. In the nonhermitian case when $\mathrm{Im}({m}(x))\neq 0$ or $\mathrm{Im}({v}(x))\neq 0$ instead, it is not possible to choose $A_{1,2}$ such that the solutions are real: In other words, the solutions are complex for all choices of $A_{1,2}$. This also mandates that the complex phase of the wavefunction is not uniform in the nonhermitian case.
@@ IV.3 Length scales and hairstyles
As we have shown, the existence and number of zero modes and their pseudospin in the infinite interval $-\infty<x<\infty$ only depend on the topological mass ${M}_{L,R}$ and topological invariants $W_{L,R}$, i.e., on the values of the fields ${m}_{L,R}$ and ${v}_{L,R}$ at large distances $|x|>w$, with their pseudospin determined only by $W_{L,R}$. These modes are localized at the smooth domain wall, and their number coincides with the difference between the topological invariants on the left and right sides $|W_{R}-W_{L}|$. Analogously, the existence of a zero mode in the semi-infinite interval $0\leq x<\infty$ only depends on the mass term ${M}_{R}$ and topological invariant $W_{R}$, i.e., on the values of the fields ${m}_{R}$ and ${v}_{R}$ at $x>w$, with its pseudospin determined only by $W_{R}$. This mode is localized at an infinite wall at $x=0$, when the topological invariant on the right side is nontrivial $|W_{R}|=1$. These results generalize the bulk-boundary correspondence hatsugai_chern_1993 ; ryu_topological_2002 ; teo_topological_2010 nonhermitian systems. On the other hand, the quantities $W_{L,R}$ in Eq. 11c generalize the topological invariants to nonhermitian systems.
@@
At large distances $x/w\to\pm\infty$, the asymptotic behavior of each mode in Eq. 8 is determined only by the associated decay rates $\mu_{L,R}^{\pm}$ and momenta $\kappa_{L,R}^{\pm}$ or, alternatively, by the characteristic decay lengths $\xi_{L,R}^{\pm}$ and oscillation wavelengths $\lambda_{L,R}^{\pm}$ for each asymptote and for each mode, describing the exponential decay or exponentially damped oscillations, which ultimately depend only on the values of the fields ${m}_{L,R}$ and ${v}_{L,R}$ at large distances $|x|\gg w$. On the left asymptote, the case ${K}_{L}=0$ mandates $\kappa_{L}^{\pm}=0$ (i.e., $\lambda_{L}^{\pm}=\infty$), which corresponds to an exponential decay (without oscillations), while the case ${K}_{L}\neq 0$ mandates $\kappa_{L}^{\pm}\neq 0$ (i.e., $\lambda_{L}^{\pm}<\infty$) which correspond to exponentially damped oscillations. If $\mathrm{Im}({m}_{L})=\mathrm{Im}({v}_{L})=0$, the case ${K}_{L}=0$ is realized for ${v}_{L}^{2}+{m}_{L}\geq 0$ (exponential decay) while the case ${K}_{L}\neq 0$ is realized for ${v}_{L}^{2}+{m}_{L}<0$ (exponentially damped oscillations). Moreover, one has that $\mu_{L}^{+}=\mu_{L}^{-}$ if $\mathrm{Re}({q}_{L})=0$, and $\kappa_{L}^{+}=-\kappa_{L}^{-}$ if $\mathrm{Im}({v}_{L})=0$. Analogous statements hold for the right asymptote.
@@
Table 2: The number of modes and pseudospin $s$ satisfying the boundary conditions $\varphi(\pm\infty)=0$ on the infinite interval $-\infty<x<\infty$ and $\varphi(0)=\varphi(\infty)=0$ on the semi-infinite interval $0\leq x<\infty$ depending on the asymptotic values ${m}_{L,R},{v}_{L,R}\in\mathbb{C}$ via the quantities ${M}_{L,R}$, ${K}_{L,R}$, and $W_{L,R}$ defined in Eqs. 11 and 11c. The modes show exponential decay if ${K}_{L,R}=0$ or exponentially damped oscillations if ${K}_{L,R}\neq 0$ for $x\to\mp\infty$. In the case ${m}_{L,R},{v}_{L,R}\in\mathbb{R}$, which gives $\operatorname{sgn}({M}_{L,R})=\operatorname{sgn}({m}_{L,R})$ and ${K}_{L,R}=|\mathrm{Im}({q}_{L,R})|$, the modes show exponential decay if ${{v}}_{L}^{2}+{m}_{L}\geq 0$ or exponentially damped oscillations if ${{v}}_{L}^{2}+{m}_{L}<0$. The asymptotic behavior of the solutions is completely determined by the decay rates and decay lengths $\mu_{L,R}^{\pm}=1/\xi_{L,R}^{\pm}=\mathrm{Re}(\alpha_{L,R}^{\pm})$ and the momenta and wavelengths $\kappa_{L,R}^{\pm}=2\pi/\lambda_{L,R}^{\pm}=\mathrm{Im}(\alpha_{L,R}^{\pm})$ responsible for oscillatory behaviors. Moreover, $\mu_{L,R}^{+}=\mu_{L,R}^{-}=\mp s\mathrm{Re}({v}_{L,R})$ if $\mathrm{Re}({q}_{L,R})=0$ and $\kappa_{L,R}^{+}=-\kappa_{L,R}^{-}=\mathrm{Im}({q}_{L,R})$ if $\mathrm{Im}({v}_{L,R})=0$. In the asymptotically hermitian case $\mathrm{Im}({m}_{L,R})=\mathrm{Im}({v}_{L,R})=0$ (i.e., hermitian at large distances), then $\kappa_{L,R}^{+}=\kappa_{L,R}^{-}=0$ for ${{v}}_{L,R}^{2}+{m}_{L,R}\geq 0$, with $\mu_{L,R}^{+}=\mu_{L,R}^{-}$ for ${{v}}_{L,R}^{2}+{m}_{L,R}=0$, while $\mu_{L,R}^{+}=\mu_{L,R}^{-}=\mp s{v}_{L,R}$ and $\kappa_{L,R}^{+}=-\kappa_{L,R}^{-}=\mathrm{Im}({q}_{L,R})\neq 0$ for ${{v}}_{L,R}^{2}+{m}_{L,R}<0$.
At short distances $|x|\lesssim w$ instead, the modes cannot be fully described only by the characteristic lengths $\xi_{L,R}^{\pm}$, $\lambda_{L,R}^{\pm}$ (or by the complex numbers $\alpha_{L,R}^{\pm}$). This is because, at short distances, the wavefunctions of the zero modes also depend on the width of the smooth domain wall $w$ and on the details of the spatial dependences of the fields. For instance, the wavefunctions of the zero modes in Eq. 7 depend on the value ${m}_{D}={m}(0)$, which is not fully determined by the asymptotic values ${m}_{L,R}$.
@@
In the case of a sharp domain wall $w\to 0$, the zero modes are only characterized by the lengths $\xi_{R}$, $\lambda_{R}$ (alternatively, by one complex number $\alpha_{R}$) on the right of the domain wall $x>0$, and by another set of characteristic lengths $\xi_{L}$, $\lambda_{L}$ (or alternatively, by one complex number $\alpha_{L}$) on the left of the domain wall $x<0$. These zero modes localized at sharp phase boundaries are featureless objects with "no hair", using the terminology introduced in Ref. marra_topological_2024, . Likewise, zero modes localized at a smooth domain wall have "hair", i.e., they are nonfeatureless, since they cannot be described entirely by a small set of numbers at all length scales. We distinguish the cases of zero modes with "short hair" when $w<\xi$ and with "long hair" when $w>\xi$. Zero modes with short hair appear featureless at large distances $|x|>\xi>w$ but not at small distances $|x|<w$, where their wavefunctions crucially depend on the detail of the fields near the domain wall. Zero modes with long hair have instead wavefunctions that are nonfeatureless at all length scales and crucially depend on the detail of the fields. Notice also that the property of having short or long hair depends on the side considered since a zero mode can have short hair on the left $x<0$ and long hair on the right of the domain wall $x>0$, or viceversa. The existence of zero modes localized at smooth domain walls is still a direct consequence of the bulk-boundary correspondence in the hermitian regime and of the generalized form of the bulk-boundary correspondence (in the language of the generalized invariant in Eq. 11c) in the nonhermitian regime.

### v2

@@ III.3.1 One zero mode for $W_{R}=\pm 1$
For $\operatorname{sgn}({M}_{R})=-1$ and $\mathrm{Re}({v}_{R})\neq 0$, the topological invariant is nonzero $W_{R}=\operatorname{sgn}(\mathrm{Re}({v}_{R}))\neq 0$, giving $\operatorname{sgn}(\mathrm{Re}(\alpha_{R}^{\pm}))=s\operatorname{sgn}(\mathrm{Re}({v}_{R}))$. Choosing $s=-\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ gives $\mathrm{Re}(\alpha_{R}^{\pm})<0$, and therefore the two linearly independent solutions $\varphi_{1,2}^{s}(x)$ in Eq.˜42 diverge for $x\to\infty$ with different decaying lengths, and thus their linear combination cannot satisfy the boundary conditions. Choosing instead $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ gives $\mathrm{Re}(\alpha_{R}^{\pm})>0$, and with both solutions $\varphi_{1,2}^{s}(x)$ decaying to zero for $x\to\infty$. Therefore, the eigenmode satisfying the boundary conditions is
@@
with pseudospin $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ and $\varphi_{1,2}^{s}(x)$ as in Eqs.˜42 and 57, $\alpha_{L}$}, $\alpha_{R}$} given by Eq.˜37, and with $A_{1,2}$ chosen such that $\varphi(0)=A_{1}\varphi_{1}^{s}(0)+A_{2}\varphi_{2}^{s}(0)=0$. Since $\mathrm{Re}(\alpha_{R}^{-})\leq\mathrm{Re}(\alpha_{R}^{+})$, the asymptotic behavior is
@@
where $\mu_{R}^{-}=\mathrm{Re}(\alpha_{R}^{-})=-{M}_{R}$. If one can expand m(x){m}(x) and v(x){v}(x) as in Eq.˜46, taking $F_{1,2}$ as in Eq.˜57, the constants $A_{1,2}$ must satisfy
@@
Notice that, when considering the eigenmodes of the modified Jackiw-Rebbi equation on the interval $0,\infty)$, the choice of the fields at $x=$-$\infty$ is somewhat arbitrary. Therefore, one can always choose the values of ${v}_{L}$ and ${m}_{L}$ in order to match the conditions ${v}_{L,R}={v}$, ${v}_{L,R}={v}$ and ${m}_{L,R}={m}$, ${v}_{L,R}=\pm{v}$, ${v}_{L,R}=\pm{v}$ and ${m}_{L,R}={m}$ in Table I of Ref. marra_topological_2024 . and obtain the respective simplified forms of the wavefunctions. However, if one can expand m(x){m}(x) and v(x){v}(x) as in Eq.˜46, the functional form of the fields m(x){m}(x) and v(x){v}(x) on the interval $0,\infty)$ determines univocally the values of the fields at $x=$-$\infty$.
The asymptotic behavior is determined by the exponent $\alpha_{R}^{-}=s{v}_{R}-{q}_{R}$ (decay length $\xi_{R}^{-}$ and oscillation wavelength $\lambda_{R}^{-}$). This is because, even though the two solutions have different decaying lengths $\xi_{R}^{\pm}$, one has that $\xi_{R}^{-}\geq\xi_{R}^{+}$, and therefore the decaying length $\xi_{R}^{-}$ dominates at large distances. In the hermitian case $\mathrm{Im}({m}(x))=\mathrm{Im}({v}(x))=0$, it is always possible to find a set of $A_{1,2}$ satisfying the equations above and such that $\varphi(x)$ is real.
@@ III.3.2 No zero modes otherwise
If $\operatorname{sgn}({M}_{R})\geq 0$ or if $\mathrm{Re}({v}_{R})=0$, no nontrivial linear combination of the general solutions can satisfy the boundary conditions. For $\operatorname{sgn}({M}_{R})=1$, the topological invariant is zero $W_{R}=0$, and one has $\mathrm{Re}(\alpha_{R}^{+})>0>\mathrm{Re}(\alpha_{R}^{-})$. Therefore only the solution $\varphi_{2}^{s}(x)$ in Eq.˜42 decays for $x\to\infty$. For $\operatorname{sgn}({M}_{R})=0$, one has that either $\alpha_{R}^{+}=0$ or $\alpha_{R}$, depending on the choice of the pseudospin. Therefore, only one of the solutions $\varphi_{1,2}^{s}(x)$ decays for $x\to\infty$. However, since in general $\varphi_{1,2}^{s}(0)\neq 0$, the boundary conditions cannot be satisfied. For ${m}_{R}<0$ (or $\operatorname{sgn}({M}_{R})=-1$) and $\mathrm{Re}({v}_{R})=0$, one has $\mathrm{Re}(\alpha_{R}^{\pm})=0$. Hence, neither solutions $\varphi_{1,2}^{s}(x)$ decay for $x\to\infty$ and thus the boundary conditions cannot be satisfied.
@@ IV.1 Oscillating behavior
Here, we will see that the quantity ${K}_{L,R}$ will discriminate between oscillatory and nonoscillatory behavior, generalizing to the nonhermitian case the analogous quantity introduced in Ref. marra_topological_2024 for hermitian systems. One has
[…]
@@
where we remind that $\mathrm{Re}(\alpha_{L}^{\pm})=\mu_{L}^{\pm}$ and $\mathrm{Im}(\alpha_{L}^{\pm})=\kappa_{L}^{\pm}$. The case where ${K}_{L}\neq 0$ corresponds to oscillating behavior for $x\to$-$\infty$, while ${K}_{L}=0$ corresponds to no oscillations. If ${m}_{L}\in\mathbb{R}$, we see that $\operatorname{sgn}({M}_{L})=\operatorname{sgn}({m}_{L})$, and thus the conditions in Eq.˜11 simplify by substituting ${M}_{L}\to{m}_{L}$ (analogously the conditions in Eq.˜12 simplify by substituting ${M}_{R}\to{m}_{R}$). Moreover, if the system is asymptotically hermitian for $x\to$-$\infty$, i.e., if ${m}_{L},{v}_{L}\in\mathbb{R}$, one has that ${K}_{L}=0$ if ${{v}}_{L}^{2}+{m}_{L}\geq 0$ and ${K}_{L}\neq 0$ otherwise. This gives
[…]
@@
The oscillatory behavior of the eigenmodes is determined by ${K}_{L,R}$. For ${K}_{L,R}=0$, one gets $\kappa_{L,R}=0$, respectively, which mandates that the modes decay exponentially for $x\to\mp\infty$, respectively. Conversely, for ${K}_{L,R}\neq 0$, one has $\kappa_{L,R}\neq 0$, respectively, which mandates modes with exponentially damped oscillations for $x\to\mp\infty$, respectively. In short, one has
@@
This is also summarized in Table˜2.
@@ IV.2 Reality
Note also that, in the hermitian case $\mathrm{Im}({m}(x))=\mathrm{Im}({v}(x))=0$, the modified Jackiw-Rebbi equation is real, and consequently, all eigenmodes are real for some choices of $A_{1,2}$. In the nonhermitian case when $\mathrm{Im}({m}(x))\neq 0$ or $\mathrm{Im}({v}(x))\neq 0$ instead, it is not possible to choose $A_{1,2}$ such that the solutions are real: In other words, the solutions are complex for all choices of $A_{1,2}$. This also mandates that the complex phase of the wavefunction is not uniform in the nonhermitian case.
@@ IV.3 Length scales and hairstyles
As we have shown, the existence and number of zero modes only depend on the topological invariants $W_{L,R}$ and ultimately on the asymptotic values of the fields ${m}_{L,R}$ and ${v}_{L,R}$. These results generalize the bulk-boundary correspondence hatsugai_chern_1993 ; ryu_topological_2002 ; teo_topological_2010 to nonhermitian systems. On the other hand, the quantities $W_{L,R}$ in Eq.˜10c generalize the topological invariants to nonhermitian systems. Moreover, the asymptotic behavior of each mode in Eq.˜7 is determined only by the decay rates $\mu_{L,R}^{\pm}$ and oscillation momenta $\kappa_{L,R}^{\pm}$ (or, alternatively, by the decay lengths $\xi_{L,R}^{\pm}$ and oscillation wavelengths $\lambda_{L,R}^{\pm}$), which ultimately depend only on the asymptotic values of the fields ${m}_{L,R}$ and ${v}_{L,R}$. On the left asymptote, the case ${K}_{L}=0$ mandates $\kappa_{L}^{\pm}=0$ (i.e., $\lambda_{L}^{\pm}=\infty$), which corresponds to an exponential decay (without oscillations), while the case ${K}_{L}\neq 0$ mandates $\kappa_{L}^{\pm}\neq 0$ (i.e., $\lambda_{L}^{\pm}<\infty$) which correspond to exponentially damped oscillations. If $\mathrm{Im}({m}_{L})=\mathrm{Im}({v}_{L})=0$, the case ${K}_{L}=0$ is realized for ${v}_{L}^{2}+{m}_{L}\geq 0$ (exponential decay) while the case ${K}_{L}\neq 0$ is realized for ${v}_{L}^{2}+{m}_{L}<0$ (exponentially damped oscillations). Moreover, one has that $\mu_{L}^{+}=\mu_{L}^{-}$ if $\mathrm{Re}({q}_{L})=0$, and $\kappa_{L}^{+}=-\kappa_{L}^{-}$ if $\mathrm{Im}({v}_{L})=0$. Analogous statements hold for the right asymptote.
infinite interval $(-\infty,\infty)$
$W_{L}W_{R}=-1$ 2 modes $\begin{cases}\varphi_{1}^{s}(x),\\ \varphi_{2}^{s}(x),\end{cases}$ $s=-W_{L}=W_{R}$
${M}_{L}{M}_{R}\neq 0$
$W_{L}=0$, $|W_{R}|=1$
${M}_{L}{M}_{R}\neq 0$
1 mode $\varphi_{1}^{s}(x)$ $s=W_{R}$
$W_{L}W_{R}=-1$ (limiting case)
${M}_{L}=0$, ${M}_{R}\neq 0$
$|W_{L}|=1$, $W_{R}=0$
${M}_{L}{M}_{R}\neq 0$
1 mode $\varphi_{2}^{s}(x)$ $s=-W_{L}$
$W_{L}W_{R}=-1$ (limiting case)
${M}_{L}\neq 0$, ${M}_{R}=0$
otherwise none
semi-infinite interval $0,\infty)$
$|W_{R}|=1$ 1 mode $A_{1}\varphi_{1}^{s}(x)+A_{2}\varphi_{2}^{s}(x)$ $s=W_{R}$
${M}_{R}\neq 0$
otherwise none
@@
Table 2: The number of eigenmodes $\varphi_{1,2}^{s}(x)$ satisfying the boundary conditions and their corresponding pseudospin $s$ on the infinite interval $(-\infty,\infty)$ and on the semi-infinite interval $0,\infty)$ for different regimes determined by the quantities ${M}_{L,R}$, ${K}_{L,R}$, and $W_{L,R}$ (topological invariant) defined in Eq.˜10, which ultimately depend on the the asymptotic values of the fields ${m}_{L,R},{v}_{L,R}\in\mathbb{C}$. The modes have exponential decay when ${K}_{L,R}=0$ or exponentially damped oscillations when ${K}_{L,R}\neq 0$ for $x\to\mp\infty$. In the asymtpotical hermitian regime (i.e., hermitian at large distances) where ${m}_{L,R},{v}_{L,R}\in\mathbb{R}$, which gives $\operatorname{sgn}({M}_{L,R})=\operatorname{sgn}({m}_{L,R})$ and ${K}_{L,R}=|\mathrm{Im}({q}_{L,R})|$, the modes have exponential decay when ${{v}}_{L}^{2}+{m}_{L}\geq 0$ or exponentially damped oscillations when ${{v}}_{L}^{2}+{m}_{L}<0$. The asymptotic behavior is fully determined by the decay rates $\mu_{L,R}^{\pm}=\mathrm{Re}(\alpha_{L,R}^{\pm})$ and the oscillation momenta $\kappa_{L,R}^{\pm}=\mathrm{Im}(\alpha_{L,R}^{\pm})$, or alternatively by the decay lengths $\xi_{L,R}^{\pm}=1/\mathrm{Re}(\alpha_{L,R}^{\pm})$ and oscillation wavelengths $\lambda_{L,R}^{\pm}=2\pi/\mathrm{Im}(\alpha_{L,R}^{\pm})$. Moreover, $\mu_{L,R}^{+}=\mu_{L,R}^{-}=\mp s\mathrm{Re}({v}_{L,R})$ if $\mathrm{Re}({q}_{L,R})=0$ and $\kappa_{L,R}^{+}=-\kappa_{L,R}^{-}=\mathrm{Im}({q}_{L,R})$ if $\mathrm{Im}({v}_{L,R})=0$. In the asymptotically hermitian regime, one has $\kappa_{L,R}^{+}=\kappa_{L,R}^{-}=0$ for ${{v}}_{L,R}^{2}+{m}_{L,R}\geq 0$, with $\mu_{L,R}^{+}=\mu_{L,R}^{-}$ for ${{v}}_{L,R}^{2}+{m}_{L,R}=0$, while $\mu_{L,R}^{+}=\mu_{L,R}^{-}=\mp s{v}_{L,R}$ and $\kappa_{L,R}^{+}=-\kappa_{L,R}^{-}=\mathrm{Im}({q}_{L,R})\neq 0$ for ${{v}}_{L,R}^{2}+{m}_{L,R}<0$. This Table summarizes the existence and number of zero-energy eigenmodes of the Jackiw-Rebbi equation in the nonhermitian case. The hermitian case can be obtained as a special case. For a direct comparison to the hermitian case, see Table II of Ref. marra_zero_2025 .
@@
In the case of a sharp domain wall $w\to 0$, the zero modes are featureless objects with "no hair" fully described by $\mu_{L,R}^{\pm}$ and $\kappa_{L,R}^{\pm}$ (or, alternatively, by $\xi_{L,R}^{\pm}$ and $\lambda_{L,R}^{\pm}$), to use the terminology introduced in Ref. marra_topological_2024 . In the case of a smooth domain wall, however, the zero modes are nonfeatureless, since they cannot be fully described by $\mu_{L,R}^{\pm}$ and $\kappa_{L,R}^{\pm}$ (or by $\xi_{L,R}^{\pm}$ and $\lambda_{L,R}^{\pm}$), especially in the region close to the domain wall $|x|<w$. We further distinguish the cases of zero modes with "short hair" when $w<\xi$ and with "long hair" when $w>\xi$. Zero modes with short hair appear featureless at large distances $|x|>\xi>w$ but not at small distances $|x|<w$, where their wavefunctions crucially depend on the detail of the fields near the domain wall. Zero modes with long hair have instead wavefunctions that are nonfeatureless at all length scales and crucially depend on the details of the fields. Notice also that the property of having short or long hair depends on the side considered since a zero mode can have short hair on the left $x<0$ and long hair on the right of the domain wall $x>0$, or vice versa. The existence of zero modes localized at smooth domain walls is still a direct consequence of the bulk-boundary correspondence in the hermitian regime and of the generalized form of the bulk-boundary correspondence (in the language of the generalized invariant in Eq.˜10c) in the nonhermitian regime.
We stress that the terminology used here is meant to be taken with a grain of salt. Black holes are said to have no hair since each black hole can be completely described (in the frame of reference of its center of mass) by only a few physical quantities, namely, its mass, electric charge, and angular momentum. Similarly, zero modes with no hair can be completely described by only a few physical quantities, namely the decay rates $\mu_{L,R}^{\pm}$ and oscillation momenta $\kappa_{L,R}^{\pm}$ (or alternatively decay lengths $\xi_{L,R}^{\pm}$ and oscillation wavelengths $\lambda_{L,R}^{\pm}$). Zero modes with hair cannot be fully described by these quantities alone. However, there is a fundamental difference between hairless zero modes and black holes: Zero modes are local objects, while black holes are not. Indeed, the wavefunctions of zero modes decay exponentially, and therefore zero modes cannot be directly observable at a distance much larger than their localization lengths. It is simply not possible to measure their decay rates and oscillation momenta at a distance larger than their localization length. In contrast, a black hole induces gravitational and electromagnetic fields which scale polynomially and not exponentially, and consequently its mass, charge, and angular momentum are observable at all distances and independent from the observer’s distance.